OpenStudy (mysangels):

Please help! A swing loses 20% of its previous height on each swing. Suppose the swing starts at a height of 10 ft. Choose the answer that represents the height of the swing after its fourth swing. A) 6.4 ft B) 5.1 ft C) 8.0 ft D) 4.1 ft

OpenStudy (narissa):

height after first swing is 10 - 0.2(10) = 8 height after second swing is 8 - 0.2(8) = 6.4 height after third swing is 6.4 - 0.2(6.4) = 5.12 height after fourth swing is 5.12 - 0.2(5.12) = 4.1 ft answer d

OpenStudy (mysangels):

How would you use the equation A(n) = 10 x 1/5 ^ n-1? This is a sequence question, but I kept getting 0.016. How does that get to 5.1 ft?